There are three hundred people at a concert.
One hundred and fifty-five of them are male. How many are female?

Answer by mathmate(429)


Question:
There are three hundred people at a concert.
One hundred and fifty-five of them are male. How many are female?

Solution:
Algebraically,
Let F=number of females
155+F=300
Transpose and solve for F:
F=300-155=145
There are 145 female.

Or, arithmetically.
The sum of males and females is 300, there are 155 males, so
number of females = 300-155=145